# Glassy word problems: ultraslow relaxation, Hilbert space jamming, and computational complexity

### 07 Dec 2023 v1

@@ Appendix B A primer in geometric group theory
A comprehensive survey of geometric properties of ${\sf EL}$ can be found in Ref. Riley (2006). Similarly to the Dehn function, the asymptotic scaling of ${\sf EL}(L)$ is independent of the choice of (finite) presentation. Also as with the Dehn function, considering expansion in other $K_{g}$ sectors does not yield anything new. Just as with proposition 4, one can similarly show that
@@
since a path in the Cayley graph can only reach $N$ different vertices before doubling back on itself. Let $M={\sf EL}(N)$. If we append $M$ identity characters to the end of $w$, we can use them to turn any one of the $w_{j}$ into $e$ without increasing the length of the appended word. Since we can do this for all of the $w_{j}$, we thus have ${\sf EL}(L)=L+N$ as claimed. ∎
Furthermore, we will see later that despite having exponential time complexity (${\sf Dehn}(L)\sim 2^{L}$), ${\sf BS}(1,2)$ only has linear spatial complexity (${\sf EL}(L)\sim L$). This is part of a more general result that all so-called asychronously combable groups, of which ${\sf BS}(1,2)$ is one, have ${\sf EL}(L)\lesssim L$ Gersten (1993). It is however relatively straightforward to construct examples where ${\sf EL}(L)$ grows faster than linearly. This is done simply by finding groups with Dehn functions that scale as ${\sf Dehn}(L)=\omega(2^{L})$. This is due to the “spacetime” bound mentioned in the main text:
Proposition 8 (Gersten and Riley (2002)).
@@
For $w\in K_{e}$ with expansion length ${\sf EL}(w)$, the number of words that $w$ can possibly visit as it is reduced to the identity is |$|\{a_{i},a_{i}^{-1},e\}|^{{\sf EL}(w)}=(2n_{g}+1)^{{\sf EL}(w)}$. Since the quickest reduction of $w$ to $e$ cannot visit a given word $w^{\prime}$ more than once, the number of steps in the reduction is (often very loosely) upper bounded by $(2n_{g}+1)^{{\sf EL}(w)}$. ∎
@@
A simple measure of a group’s geometry is how fast the group grows, viz. how the number of elements within a distance $L$ of the origin of the Cayley graph grows as $L$ is increased. To this end, let $\mathcal{B}(L)$ denote the radius-$L$ ball centered at the origin of the Cayley graph, viz.
@@
and define the growth function (or group volume) as $N_{K}(L)\triangleq|\mathcal{B}(L)|$. Physically, $N_{K}(L)$ places a lower bound on the number of Krylov sectors that models with ${\sf Dyn}_{G}$ dynamics fragment into ($N_{K}(L)$ equalling the number of sectors in the absence of fragile fragmentation). A basic property of a group is the asymptotic scaling of $N_{K}$ with $L$. It is straightforward to $N_{K}\sim O(1)$ for finite groups and $N_{K}\sim{\rm poly}(L)$ for Abelian groups. Free groups provide the simplest examples where $N_{K}\sim\exp(L)$, which is the maximum possible growth rate. [^31] Groups with growth rate intermediate between polynomial and exponential exist — the “simplest” examples are groups arising from automorphisms of trees Grigorchuk and Pak (2008) — but it is conjectured that all such groups have only infinite presentations. We now ask what implications group expansion has for the time and space complexity measures introduced above. It turns out that exponential growth is needed in order to have ${\sf Dehn}(L)>{\rm poly}(L)$:
@@
The contrapositive of this proposition follows from Gromov’s aforementioned result that all finitely generated groups with polynomial growth are virtually nilpotent (viz. have a nilpotent subgroup of finite index) Gromov (1981). Since virtually nilpotent groups have the same Dehn functions as nilpotent ones, we may combine Gromov’s theorem with the fact that all nilpotent groups have ${\sf Dehn}(L)\sim L^{d}$ for some $d$ Gersten et al. (2003) to arrive at the result. ∎
@@
where $\Delta$ is the normalized graph Laplacian and $\delta_{L}$ denotes the discrete derivative along the “time” direction determined by $L$. Thus the different sizes of the $K_{g,L}$ are determined by using the heat equation to evolve a delta function concentrated on $e$ for a total time of $L$. The claim we are trying to prove then follows from estimates of the discrete heat kernel Greens function developed in the graph theory literature, see e.g. Davies (1992) for a review. ∎
@@
We aim to show that $p_{L}(e,e)\geq p_{L}(e,g)$ for all $g,L$. For simplicity of notation, let $L\in 2\mathbb{N}$. Then using $p_{L}(g,h)=p_{L}(gk,hk)$ for all $g,h,k$ on account of translation invariance of the Cayley graph, we have
@@
In the main text, we saw numerically that this bound is saturated for ${\sf BS}$ (see Fig. 4). This means that ${\sf Dyn}_{\sf BS}$ is as close to being weakly fragmented as a group dynamics model with exponentially slow relaxation can be. We can also connect the above bound with a previous result to derive:
@@ Appendix C Semigroup dynamics and undecidability
While not relevant to the examples studied in the main text, it is amusing to note that the computational problems of relevance to semigroup dynamics cannot only be computationally hard, but can even be undecidable. For example, a foundational result in computability theory is the undecidability of the semigroup word problem, as first proven by Markov Markov (1948). This result implies the existence of finitely-presented semigroups for which ${\sf Dehn}(w,w^{\prime})$ grows faster than any recursive function. For such groups, one can never be sure whether or not two words $|w\rangle,|w^{\prime}\rangle$ are in the same $K_{g}$, showing that establishing the number of dynamical sectors is impossible in general. Remarkably, rather simple examples of semigroups with undecidable word problems are known, with the required $|\mathcal{H}_{\rm loc}|$ being as low as five. An explicit example from Ref. Tseitin (1958) serves as an illustration how simple the examples can get: the semigroup in question has a five-element generating set $S=\{\mathtt{v},\mathtt{w},\mathtt{x},\mathtt{y},\mathtt{z}\}$, and the seven relations
@@
The word problem is also undecidable when one specifies further to the setting of a finitely presented group Novikov (1955); Boone (1959). Going further still, the Adyan-Rabin theorem Rabin (1958); Adyan (1955) states that nearly all “reasonable” properties of finitely-presented groups—their Dehn times, whether or not they are finite, Abelian; even whether or not they are a presentation of the trivial group—is undecidable. As a corollary, the existence of Hilbert space fragmentation itself is therefore undecidable: even if ergodicity looks to be broken for all system sizes below $L$, for some group presentations one can in general never be sure that it will not be restored at system size $L+1$. This adds on to recent work on the undecidability of physical problems relating to local Hamiltonians, e.g. the determination of spectral gaps, ground state phase diagrams, and so on Cubitt et al. (2015); Bausch et al. (2020); Shiraishi and Matsumoto (2021); Bausch et al. (2021).
@@ Appendix D No complete symmetry labels for non-Abelian $G$
We will argue by contradiction. Assume that there existed a set of unitaries $\mathcal{U}_{a}$ whose expectation values in a given computational basis product state $|w\rangle$ allowed one to determine the group element $\varphi(w)$ associated with $w$ (and therefore the $K_{g,L}$ that $|w\rangle$ belongs to). Consider the states
@@
Since we have assumed $m$ to be extensive and $\mathcal{U}_{a}$ to be locality-preserving, we will always have $m+1>2\,{\rm supp}(\mathcal{O}^{\mathcal{U}_{a}}_{\mathtt{g},i})$ in the thermodynamic limit, and so the above expectation value splits as
@@
Taking $\mathtt{g},\mathtt{h}$ to be any two generators such that $\mathtt{g}\mathtt{h}\neq\mathtt{h}\mathtt{g}$ as group elements, the above then shows that states in different $K_{g,L}$ cannot be distinguished by eigenvalues of the $\{\mathcal{U}_{a}\}$, provided that i) the above assumption about translation invariance can be removed, and ii) $|w_{\mathtt{g}\mathtt{h}}\rangle$ is not orthogonal to $|w_{\mathtt{g}\mathtt{h};a}\rangle$ for all choices of $l,m,$ and all choices of $\mathtt{g},\mathtt{h}$ such that $\mathtt{g}\mathtt{h}\neq\mathtt{h}\mathtt{g}$. i) is dealt with by noting that when the $\mathcal{U}_{a}$ are not translation invariant (as is the case for e.g. modulated symmetries), we may exploit the fact that the sectors to which $|w_{\mathtt{g}\mathtt{h}}\rangle$ and $|w_{\mathtt{h}\mathtt{g}}\rangle$ belong are independent of $l,m$, and hence $l,m$ can be varied without changing the eigenvalues of the respective states under the $\{\mathcal{U}_{a}\}$. For ii), we need only note that
@@
which is always nonzero by the unitarity of $\mathcal{U}_{a}$. ∎
@@ Appendix E The geometry and complexity of Baumslag-Solitar groups
In this section we state and prove some facts about the group geometry of the Baumslag-Solitar group ${\sf BS}(1,2)$ and some of its simple generalizations. The original work in which these groups were defined is Ref. Baumslag and Solitar (1962). We will always work with the presentation
@@
Everything we say below can in fact be modified only slightly in order to work for the groups ${\sf BS}(1,q)=\langle\mathtt{a},\mathtt{b}\,|\,\mathtt{a}\mathtt{b}=\mathtt{b}\mathtt{a}^{q}\rangle$, but for concreteness we will specify to $q=2$ throughout, and as in the main text will write ${\sf BS}$ for ${\sf BS}(1,2)$. A useful fact about our chosen presentation for ${\sf BS}$ is that it admits the the matrix representation Clay and Margalit (2017)
@@
We will start by analyzing the worst-case time and space complexity of words in ${\sf BS}$, which we do by computing the Dehn and expansion length functions. To orient ourselves, it is helpful to recall that the Cayley graph of ${\sf BS}$ is homeomorphic to the product of the real line with a 3-regular tree (for ${\sf BS}(1,q)$ it is a $(q+1)$-tree), with the depth of a given word $w$ along the tree being controlled by the relative number of $b$s and $b^{-1}$s that $w$ contains (see Fig. 3 of the main text). We will refer to this tree as the b-tree, and will use terminology whereby multiplying by $b$ ($b^{-1}$) moves one “up” (“down”) on the sheet of the $b$-tree one is currently at, while multiplying by $a$ ($a^{-1}$) moves one “right” (“left”) within the sheet (along the a-axis). Because of the hierarchical nature of the graph, multiplying by $b$ — viz. moving deeper into the $b$-tree — moves one to “larger scales”, while multiplying by $b^{-1}$ does the opposite.
@@
We now turn to computing the Dehn function and expansion length of ${\sf BS}$. The Dehn function must be large, since it is easy to construct words with ${\sf Dehn}(w)\sim 2^{|w|}$ Gersten (1992):
@@
A visual illustration of this statement is given in Fig. 3. The proof, which is a condensed version of the original proof in Ref. Gersten (1992), is as follows:
@@
$22\sum_{k=0}^{n}2^{k}=2^{n+1}-$ 2 . (155)
Therefore we need only show that this area cannot be reduced by considering surfaces that extend into other sheets. This is however clearly true, as a 2-cell of $S_{w_{big}}$ that lives on any other sheet will necessarily make an unwanted contribution to $\partial S_{w_{big}}$. More formally, we can recognize that, being homeomorphic to the product of $\mathbb{R}$ with a 3-tree, the Cayley 2-complex is contractible, and thus the $S_{big}$ found above is the unique bounding surface. ∎
@@
[Gersten (1993)] ${\sf BS}$ has linear expansion length, ${\sf EL}(L)\sim L$.
The previous proposition immediately implies the existence of exponentially many (in $L$) elements of $K_{e,L}$ with exponentially large (also in $L$) area. The Dehn function thus grows at least as ${\sf Dehn}(L)\gtrsim 2^{L}$. We now show that this bound is in fact tight. The existing proof of this fact in the mathematical literature appears to be to note that ${\sf BS}$ is an “asychronously automatic” group Epstein (1992), and that any such group has ${\sf Dehn}(L)\lesssim 2^{L}$ Baumslag et al. (1991). In the following, we give a more elementary proof:
@@
The construction above tells us that ${\sf Dehn}(L)\gtrsim 2^{L}$. A matching upper bound can be proven using an algorithm which converts an input word to a particular standard form. This is done by simply moving all occurances of $b$ in $w$ to the left and all occurances of $b^{-1}$ to the right, dupicating $a$ and $a^{-1}$s along the way as needed and (optionally, for our present purposes) eliminating $bb^{-1}$ pairs as they are encountered. It is clear that at most an exponential in $L$ number of additional $a$s are generated during this process of shuffling the $b$s and $b^{-1}$s around (as usual, this means exponential up to a polynomial factor, here linear in $L$), so that the resulting word is
$w^{\prime}=b^{k}w_{a}b^{-l},$ (157)
where $w_{a}$ is a word containing only $\mathtt{e}$, $\mathtt{a},\mathtt{a}^{-1}$ and with length $|w_{a}|\lesssim 2^{L}$. Since $w\sim e$, we know that $k=l$ and that $w_{a}\sim\mathtt{e}$. Thus an additional $2^{L}$ relations suffice to reduce $w_{a}$ to $\mathtt{e}$ and hence $w^{\prime}$ to $\mathtt{e}$, and hence $2^{L}$ is also an upper bound on ${\sf Dehn}(L)$. ∎
We saw in Fact 12 that ${\sf BS}$ has linear expansion length, meaning that there exists constants $C,D$ such that ${\sf EL}(w)\leq CL+D$ for all $w\in K_{e,L}$. The argument in Ref. Gersten (1993) however does not tell us whether $C=1$ or $C>1$, which is need for determining whether or not ${\sf BS}$ exhibits fragile fragmentation. The following proposition answers this question in the affirmative:
@@
It is enough to show that ${\sf EL}(w)\geq(1+\alpha)L+D$ for a particular length-$L$ word $w$, which we choose to be $w_{big}$ of (153) with $n$ such that $L=4(n+1)$. By the contractibility of the Cayley 2-complex ${\mathsf{CG}}_{\sf BS}$, for all $0\leq m\leq 2^{n}$, any homotopy from $w_{big}$ to the identity must pass through the vertex $\mathtt{a}^{m}$ of ${\mathsf{CG}}_{\sf BS}$. Therefore
@@
where $|\mathtt{a}^{m}|$ is the geodesic distance (not word length) of $\mathtt{a}^{m}$. We now argue that there exists an $2^{n-1}<m<2^{n}$, an $\alpha>0$, and a constant $c_{1}$ such that $|\mathtt{a}^{m}|>(2+\alpha)n+c_{1}$. By the contractibility of ${\mathsf{CG}}_{\sf BS}$, any geodesic of $\mathtt{a}^{m}$ will be contained within a single sheet. Furthermore, since $m$ can be chosen to be exponentially large in $n$, the geodesic must reach a height of at least $n-c_{2}$ on the sheet, where $c_{2}$ is $O(1)$. Letting $n_{|\mathtt{b}|}$ ($n_{|\mathtt{a}|}$) be the number of $\mathtt{b},\mathtt{b}^{-1}$ ($\mathtt{a},\mathtt{a}^{-1}$) characters that appear in the geodesic, this shows that $n_{|\mathtt{b}|}\geq 2(n-c_{2})$, and so $\alpha\geq 0$. The value of $\alpha$ is then determined by $n_{|\mathtt{a}|}$. Since there are exponentially many (in $n$) choices of $m$, and since each $\mathtt{a}^{m}$ has a unique geodesic, most choices of $m$ must yield a geodesic with $n_{|\mathtt{a}|}$ proportional to $n$. In fact for a given $m$, one sees that $n_{|\mathtt{a}|}$ is determined by the runtime of a binary search algorithm applied to $m$, which is maximized for
$m=2^{n}-2^{n-1}+2^{n-2}-\cdots=2^{n-1}+2^{n-3}+\cdots$ (160)
In this case $n_{|\mathtt{a}|}=n+c_{3}$ for yet another constant $c_{3}$, and so there is a $c_{4}$ such that $\max_{0\leq m\leq 2^{n}}2|\mathtt{a}^{m}|=6n+c_{4}$, meaning that (158) is satisfied with $\alpha=1/2$. ∎

### 11 Apr 2024 v2

@@ Appendix B A primer in geometric group theory
A comprehensive survey of geometric properties of ${\sf EL}$ can be found in Ref. [70]. Similarly to the Dehn function, the asymptotic scaling of ${\sf EL}(L)$ is independent of the choice of (finite) presentation. Also as with the Dehn function, considering expansion in other $K_{g}$ sectors does not yield anything new. Just as with proposition 4, one can similarly show that
@@
since a path in the Cayley graph can only reach at most $N$ different vertices before returning to its starting point. Let $M={\sf EL}(N)$. If we append $M$ identity characters to the end of $w$, we can use them to turn any one of the $w_{j}$ into $e$ without increasing the length of the appended word. Since we can do this for all of the $w_{j}$, we thus have ${\sf EL}(L)=L+M$ as claimed. ∎
Furthermore, we will see later that despite having exponential time complexity (${\sf Dehn}(L)\sim 2^{L}$), ${\sf BS}(1,2)$ only has linear spatial complexity (${\sf EL}(L)\sim L$). This is part of a more general result that asychronously combable groups (which ${\sf BS}(1,2)$ is an example of) have ${\sf EL}(L)\lesssim L$ [71]. It is however relatively straightforward to construct examples where ${\sf EL}(L)$ grows faster than linearly. This is done simply by finding groups with Dehn functions that scale as ${\sf Dehn}(L)=\omega(2^{L})$. This is due to the “spacetime” bound mentioned in the main text:
Proposition 8 ([57]).
@@
For $w\in K_{e}$ with expansion length ${\sf EL}(w)$, the number of words that $w$ can possibly visit as it is reduced to the identity is |$|\{a_{i},a_{i}^{-1},e\}|^{{\sf EL}(w)}=(2n_{g}+1)^{{\sf EL}(w)}$. Since the shortest reduction of $w$ to $e$ cannot visit a given word $w^{\prime}$ more than once, the number of steps in the reduction is (often very loosely) upper bounded by $(2n_{g}+1)^{{\sf EL}(w)}$. ∎
@@
A simple measure of a group’s geometry is how fast the group grows, namely how the number of elements within a distance $L$ of the origin of the Cayley graph grows as $L$ is increased. To this end, let $\mathcal{B}(L)$ denote the radius-$L$ ball centered at the origin of the Cayley graph, namely
@@
and define the growth function (or group volume) as $N_{K}(L)\triangleq|\mathcal{B}(L)|$. Physically, $N_{K}(L)$ places a lower bound on the number of Krylov sectors that ${\sf Dyn}_{G}$ possesses (in the absence of fragile fragmentation, the number of Krylov sectors equals $N_{K}(L)$). A basic property of a group is the asymptotic scaling of $N_{K}$ with $L$. It is straightforward to $N_{K}\sim O(1)$ for finite groups and $N_{K}\sim{\rm poly}(L)$ for Abelian groups. Free groups provide the simplest examples where $N_{K}\sim\exp(L)$, which is the maximum possible growth rate. [^32] Groups with growth rate intermediate between polynomial and exponential exist — the “simplest” examples are groups arising from automorphisms of trees [72] — but it is conjectured that all such groups have only infinite presentations. We now ask what implications group expansion has for the time and space complexity measures introduced above. It turns out that exponential growth is needed in order to have ${\sf Dehn}(L)$ growing faster than ${\rm poly}(L)$:
@@
The contrapositive of this proposition follows from Gromov’s theorem that all finitely generated groups with polynomial growth are virtually nilpotent (namely have a nilpotent subgroup of finite index) [73]. Since virtually nilpotent groups have the same Dehn functions as nilpotent ones, we may combine Gromov’s theorem with the fact that all nilpotent groups have ${\sf Dehn}(L)\sim L^{d}$ for some $d$ [42] to arrive at the result. ∎
@@
where $\Delta$ is the normalized graph Laplacian and $\delta_{L}$ denotes the discrete derivative along the “time” direction determined by $L$. Thus the different sizes of the $K_{g,L}$ are determined by using the heat equation to evolve a delta function concentrated on $e$ for a total time of $L$. The claim we are trying to prove then follows from estimates of the discrete heat kernel Greens function developed in the graph theory literature, see [74] for a review. ∎
@@
We aim to show that $p_{L}(e,e)\geq p_{L}(e,g)$ for all $g,L$. For simplicity of notation, let $L\in 2\mathbb{N}$. Then using $p_{L}(g,h)=p_{L}(gk,hk)$ for all $g,h,k$ on account of transitivity of the Cayley graph, we have
@@
In the main text, we saw numerically that this bound is saturated for ${\sf BS}$ (see Fig. 4). This means that while ${\sf Dyn}_{\sf BS}$ has exponentially slow relaxation, it is as ‘close’ to being weakly fragmented as possible. We can also connect the above bound with a previous result to derive:
@@ Appendix C Semigroup dynamics and undecidability
While not relevant to the examples studied in the main text, it is amusing to note that the word problem for semigroups is not just computationally hard, but can even be undecidable. For example, a foundational result in computability theory is the undecidability of the semigroup word problem, as first proven by Markov [76]. This result implies the existence of finitely-presented semigroups for which ${\sf Dehn}(w,w^{\prime})$ grows faster than any recursive function. For such groups, one can never be sure whether or not two words $|w\rangle,|w^{\prime}\rangle$ are in the same $K_{g}$, showing that establishing the number of dynamical sectors is impossible in general. Remarkably, rather simple examples of semigroups with undecidable word problems are known, with the required $|\mathcal{H}_{\rm loc}|$ being as low as five. An explicit example from Ref. [77] provides a rather simple example of a semigroup with an undecidable word problem on the five-element generating set $S=\{\mathtt{v},\mathtt{w},\mathtt{x},\mathtt{y},\mathtt{z}\}$, and the seven relations
@@
The word problem is also undecidable when one specifies further to the setting of a finitely presented group [78, 79]. Going further still, the Adyan-Rabin theorem [80, 81] states that nearly all “reasonable” properties of finitely-presented groups—their Dehn times, whether or not they are finite or Abelian; even whether or not they are a presentation of the trivial group—is undecidable. As a corollary, the existence of Hilbert space fragmentation itself is therefore undecidable: even if ergodicity looks to be broken for all system sizes below $L$, for some group presentations one can in general never be sure that it will not be restored at system size $L+1$. This result compliments recent work on the undecidability of physical problems relating to local Hamiltonians, e.g. computing spectral gaps, ground state phase diagrams, and so on [82, 83, 84, 85].
@@ Appendix D No complete symmetry labels for non-Abelian $G$
We will argue by contradiction. Assume that there existed a set of unitaries $\mathcal{U}_{a}$ whose expectation values in a given computational basis product state $|w\rangle$ allowed one to determine the group element $\varphi(|w\rangle)$ associated with $w$ (and therefore the $K_{g,L}$ that $|w\rangle$ belongs to). Consider the states
@@
Since we have assumed $m$ to be extensive and $\mathcal{U}_{a}$ to be locality-preserving, we will always have $m+1>2\,{\rm supp}(\mathcal{O}^{\mathcal{U}_{a}}_{\mathtt{g},i})$ for sufficiently large system sizes, and so the above expectation value splits as
@@
Taking $\mathtt{g},\mathtt{h}$ to be any two generators such that $\mathtt{g}\mathtt{h}\neq\mathtt{h}\mathtt{g}$ as group elements, the above then shows that states in different $K_{g,L}$ cannot be distinguished by eigenvalues of the $\{\mathcal{U}_{a}\}$, provided that i) the above assumption about translation invariance can be removed, and ii) $|w_{\mathtt{g}\mathtt{h}}\rangle$ is not orthogonal to $|w_{\mathtt{g}\mathtt{h};a}\rangle$ for all choices of $l,m,$ and all choices of $\mathtt{g},\mathtt{h}$ such that $\mathtt{g}\mathtt{h}\neq\mathtt{h}\mathtt{g}$. i) is dealt with by noting that when the $\mathcal{U}_{a}$ are not translation invariant (as is the case for modulated symmetries), we may exploit the fact that the sectors to which $|w_{\mathtt{g}\mathtt{h}}\rangle$ and $|w_{\mathtt{h}\mathtt{g}}\rangle$ belong are independent of $l,m$, and hence $l,m$ can be varied without changing the eigenvalues of the respective states under the $\{\mathcal{U}_{a}\}$. For ii), we need only note that
@@
which is nonzero since |$\ket{\mathtt{e}}$ is an eigenstate of $\mathcal{U}_{a}^{\dagger}$ by assumption (on account of its definite symmetry charge), and since unitarity guarantees that $\mathcal{U}_{a}^{\dagger}$ has no zero eigenvalues. ∎
@@ Appendix E The geometry and complexity of Baumslag-Solitar groups
In this section we state and prove some facts about the group geometry of the Baumslag-Solitar group ${\sf BS}(1,2)$ and some of its simple generalizations. The original work in which these groups were defined is Ref. [68]. We will always work with the presentation
@@
Everything we say below can be readily modified to work for the generalixed groups ${\sf BS}(1,q)=\langle\mathtt{a},\mathtt{b}\,|\,\mathtt{a}\mathtt{b}=\mathtt{b}\mathtt{a}^{q}\rangle$, but for concreteness we will specify to $q=2$ throughout, and as in the main text will write ${\sf BS}$ for ${\sf BS}(1,2)$. A useful fact about our chosen presentation for ${\sf BS}$ is that it admits the the matrix representation [41]
@@
We will start by analyzing the worst-case time and space complexity of words in ${\sf BS}$, which we do by computing the Dehn and expansion length functions. To orient ourselves, it is helpful to recall that the Cayley graph of ${\sf BS}$ is homeomorphic to the product of the real line with a 3-regular tree (for ${\sf BS}(1,q)$ it is a $(q+1)$-tree), with the depth of a given word $w$ along the tree being controlled by the relative number of $b$s and $b^{-1}$s that $w$ contains (see Fig. 3 of the main text). We will refer to this tree as the b-tree, and will use terminology whereby multiplying by $b$ ($b^{-1}$) moves one “up” (“down”) on the sheet of the $b$-tree one is currently at, while multiplying by $a$ ($a^{-1}$) moves one “right” (“left”) within the sheet (along the a-axis). Because of the hierarchical nature of the graph, multiplying by $b$ — namely moving deeper into the $b$-tree — moves one to “larger scales”, while multiplying by $b^{-1}$ does the opposite.
@@
We now turn to computing the Dehn function and expansion length of ${\sf BS}$. The Dehn function must be large, since it is easy to construct words with ${\sf Dehn}(w)\sim 2^{|w|}$ [56]:
@@
A visual illustration of this statement is given in Fig. 3. The proof, which is a condensed version of the original proof in Ref. [56], is as follows:
@@
$22\sum_{k=0}^{n-1}2^{k}=2^{n+1}-$ 2 . (161)
Therefore we need only show that this area cannot be reduced by considering surfaces that extend into other sheets. The reader can verify that this is true, since a 2-cell of $S_{w_{big}}$ that lives on any other sheet will necessarily make an unwanted contribution to $\partial S_{w_{big}}$. More formally, we can recognize that, being homeomorphic to the product of $\mathbb{R}$ with a 3-tree, the Cayley 2-complex is contractible, and thus the $S_{big}$ found above is the unique minimal bounding surface. ∎
@@
[[71]] ${\sf BS}$ has linear expansion length, ${\sf EL}(L)\sim L$.
The previous proposition immediately implies the existence of exponentially many (in $L$) elements of $K_{e,L}$ with exponentially large (also in $L$) area. The Dehn function thus grows at least as ${\sf Dehn}(L)\gtrsim 2^{L}$. We now show that this bound is in fact tight. The existing proof of this fact in the mathematical literature appears to be to note that ${\sf BS}$ is an “asychronously automatic” group [86], and that any such group has ${\sf Dehn}(L)\lesssim 2^{L}$ [87]. In the following, we give a more elementary proof:
@@
The construction above tells us that ${\sf Dehn}(L)\gtrsim 2^{L}$. A matching upper bound can be proven using an algorithm which converts an input word to a particular standard form. This is done by simply moving all occurrences of $b$ in $w$ to the left and all occurrences of $b^{-1}$ to the right, duplicating $a$ and $a^{-1}$s along the way as needed and (optionally, for our present purposes) eliminating $bb^{-1}$ pairs as they are encountered. It is clear that at most an exponential in $L$ number of additional $a$s are generated during this process of shuffling the $\mathtt{b}$s and $\mathtt{b}^{-1}$s around (as usual, this means exponential up to a polynomial factor, here linear in $L$), so that the resulting word is
$w^{\prime}=\mathtt{b}^{k}w_{a}\mathtt{b}^{-l},$ (163)
where $w_{a}$ is a word containing only $\mathtt{e}$, $\mathtt{a},\mathtt{a}^{-1}$ and with length $|w_{a}|\lesssim 2^{L}$. Since $w\sim e$, we know that $k=l$ and that $w_{a}\sim\mathtt{e}$. Thus an additional $2^{L}$ relations suffice to reduce $w_{a}$ to $\mathtt{e}$ and hence $w^{\prime}$ to $\mathtt{e}$ – thus, $2^{L}$ is also an upper bound on ${\sf Dehn}(L)$. ∎
We saw in Fact 12 that ${\sf BS}$ has linear expansion length, meaning that there exists constants $C,D$ such that ${\sf EL}(w)\leq CL+D$ for all $w\in K_{e,L}$. The argument in Ref. [71] however does not tell us whether $C=1$ or $C>1$, which is need for determining whether or not ${\sf BS}$ exhibits fragile fragmentation. The following proposition answers this question in the affirmative:
@@
It is enough to show that ${\sf EL}(w)\geq(1+\alpha)L+D$ for a particular length-$L$ word $w$, which we choose to be $w_{big}$ in (159) with $n$ such that $L=4(n+1)$. By the contractibility of the Cayley 2-complex ${\mathsf{CG}}_{\sf BS}$, for all $0\leq m\leq 2^{n}$, any homotopy from $w_{big}$ to the identity must pass through the vertex $\mathtt{a}^{m}$ of ${\mathsf{CG}}_{\sf BS}$. Therefore
@@
where $|\mathtt{a}^{m}|$ is the geodesic distance (not word length) of $\mathtt{a}^{m}$. We now argue that there exists an $2^{n-1}<m<2^{n}$, an $\alpha>0$, and a constant $c_{1}$ such that $|\mathtt{a}^{m}|>(2+2\alpha)n+c_{1}$. By the contractibility of ${\mathsf{CG}}_{\sf BS}$, any geodesic of $\mathtt{a}^{m}$ will be contained within a single sheet. Furthermore, since $m$ is exponentially large in $n$ in the worst case, the geodesic should reach a height of at least $n-c_{2}$ on the sheet, where $c_{2}$ is a sufficiently large $O(1)$ constant. If the geodesic does not reach such a height, then it must make a much larger number of steps in the $\mathtt{a}$ direction resulting in a larger perimeter. Letting $n_{|\mathtt{b}|}$ ($n_{|\mathtt{a}|}$) be the number of $\mathtt{b},\mathtt{b}^{-1}$ ($\mathtt{a},\mathtt{a}^{-1}$) characters that appear in the geodesic, this shows that $n_{|\mathtt{b}|}\geq 2(n-c_{2})$, and so $\alpha\geq 0$. To compute the perimeter, we have to determine the number of steps the smallest length word makes in the $\mathtt{a}$ direction in order to reach $\mathtt{a}^{m}$. Since at its highest point $n_{|\mathtt{b}|}/2$ steps down along the $\mathtt{b}$ direction are needed, we can intersperse any of these $n_{|\mathtt{b}|}/2\geq n-c_{2}$ steps with at least one step along the $\mathtt{a}$ direction. Note that if two consecutive steps along the $\mathtt{a}$ direction are taken, these can be pulled past the previous $\mathtt{b}$ step to form a single $\mathtt{a}$ step, thereby reducing the total length of the loop. Therefore, the minimal length loop has $\mathtt{b}$ steps interspersed with at most a single $\mathtt{a}$ step. Since there are exponentially many (in $n$) choices of $m$, and since each $\mathtt{a}^{m}$ has a unique geodesic (i.e. minimal sequence of $\mathtt{a}$ and $\mathtt{b}$ steps), a simple application of the pigeonhole principle guarantees that $\exists m$ such that $\geq 0.98n$ steps are along $\mathtt{a}$. This can be verified because the total number of placements of less than $0.98n$ steps along $\mathtt{a}$ in a sequence of $n_{|\mathtt{b}|}\geq(n-c_{2})$ steps along $\mathtt{b}$ is $\ll 2^{n}$. Therefore, there is a $c_{4}$ such that $\max_{0\leq m\leq 2^{n}}2|\mathtt{a}^{m}|\geq 2(2+0.98)n+c_{4}=5.96n+c_{4}$, meaning that (164) is satisfied with $\alpha=0.49$. ∎